# EvoFSM: Controllable Self-Evolution for Deep Research with Finite State Machines

### v1

@@ 1 Introduction
To address the limitations of static workflows, the research community has pivoted towards “Self-Evolution” mechanisms (Deppen et al., 2024; Wang et al., 2024; Li et al., 2024), aiming to empower agents to dynamically adjust strategies based on task feedback. However, existing self-evolution methods often employ “unconstrained rewriting,” where a Meta-Agent is granted the freedom to rewrite the entire system prompt or tools of worker agents (Schmidhuber et al., 2024). This unconstrained approach poses significant stability challenges, as agents may suffer from core instruction drift, hallucination, or the corruption of robust functional modules during self-modification, potentially causing the system to deteriorate rather than improve (Shao et al., 2024), as shown in Figure 1.
@@ 2.2 Self-Evolving Agents
Recent work has explored self-evolving agents that improve during deployment (Shinn et al., 2023; Madaan et al., 2023; Yang et al., 2023). In code and tool domains, LIVE-SWE-AGENT (Deppen et al., 2024) and STELLA (Wang et al., 2024) allow agents to synthesize tools or update their scaffolding on the fly, while ShieldLearner (Ni et al., 2025) learns reusable defenses against jailbreaks. Self-evolution has also been extended to multi-agent settings, e.g., MorphAgent (Li et al., 2024) adapts agent profiles and CoMAS (Wu et al., 2025) drives decentralized co-evolution via interaction rewards. More radical proposals such as the Huxley-Gödel Machine (Schmidhuber et al., 2024) pursue open-ended self-improvement through recursive modifications, and SE-Agent (Lin et al., 2025) optimizes reasoning trajectories through multi-step interaction. Despite these advances, many approaches rely on unconstrained prompt rewriting or global architecture search with limited structural guardrails, which can lead to instability and failed evolution (Shao et al., 2024). In contrast, our EvoFSM introduces explicit structural guardrails and experience-guided evolution to keep self-improvement stable. It adapts its strategy to each query while preserving core functionality, mitigating the instruction drift and instability commonly observed in unconstrained self-evolution.
@@ 3.1 FSM Initialization
While recent works have explored self-evolution to improve adaptability, these approaches often rely on an unconstrained rewriting mechanism (Shao et al., 2024), which frequently leads to hallucination and instruction drift. To address the uncontrollable evolution, we model the deep research process as a deterministic yet dynamic FSM. Fundamentally, the FSMs serve as a directed behavioral graph, where nodes represent specialized action phases (e.g., Search, Browse) and edges define the logical transitions between them based on runtime context. This graph-based structure grounds the agent’s behaviors, serving as a robust backbone that enables the precise, targeted self-evolving optimization described in Section 3.2. We formalize this graph-based system as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{T},\mathcal{I},\mathcal{C}\rangle$, structured to align with the dual optimization dimensions of Flow and Skill:

### v2

@@ 1 Introduction
To address the limitations of static workflows, the research community has pivoted towards “Self-Evolution” mechanisms (Deppen et al., 2024; Wang et al., 2024; Li et al., 2024), aiming to empower agents to dynamically adjust strategies based on task feedback. However, existing self-evolution methods often employ “unconstrained rewriting,” where a Meta-Agent is granted the freedom to rewrite the entire system prompt or tools of worker agents (Schmidhuber et al., 2024). This unconstrained approach poses significant stability challenges, as agents may suffer from core instruction drift, hallucination, or the corruption of robust functional modules during self-modification, potentially causing the system to deteriorate rather than improve (Shao et al., 2025), as shown in Figure 1.
@@ 2.2 Self-Evolving Agents
Recent work has explored self-evolving agents that improve during deployment (Shinn et al., 2023; Madaan et al., 2023; Yang et al., 2023). In code and tool domains, LIVE-SWE-AGENT (Deppen et al., 2024) and STELLA (Wang et al., 2024) allow agents to synthesize tools or update their scaffolding on the fly, while ShieldLearner (Ni et al., 2025) learns reusable defenses against jailbreaks. Self-evolution has also been extended to multi-agent settings, e.g., MorphAgent (Li et al., 2024) adapts agent profiles and CoMAS (Wu et al., 2025) drives decentralized co-evolution via interaction rewards. More radical proposals such as the Huxley-Gödel Machine (Schmidhuber et al., 2024) pursue open-ended self-improvement through recursive modifications, and SE-Agent (Lin et al., 2025) optimizes reasoning trajectories through multi-step interaction. Despite these advances, many approaches rely on unconstrained prompt rewriting or global architecture search with limited structural guardrails, which can lead to instability and failed evolution (Shao et al., 2025). In contrast, our EvoFSM introduces explicit structural guardrails and experience-guided evolution to keep self-improvement stable. It adapts its strategy to each query while preserving core functionality, mitigating the instruction drift and instability commonly observed in unconstrained self-evolution.
@@ 3.1 FSM Initialization
While recent works have explored self-evolution to improve adaptability, these approaches often rely on an unconstrained rewriting mechanism (Shao et al., 2025), which frequently leads to hallucination and instruction drift. To address the uncontrollable evolution, we model the deep research process as a deterministic yet dynamic FSM. Fundamentally, the FSMs serve as a directed behavioral graph, where nodes represent specialized action phases (e.g., Search, Browse) and edges define the logical transitions between them based on runtime context. This graph-based structure grounds the agent’s behaviors, serving as a robust backbone that enables the precise, targeted self-evolving optimization described in Section 3.2. We formalize this graph-based system as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{T},\mathcal{I},\mathcal{C}\rangle$, structured to align with the dual optimization dimensions of Flow and Skill: